# Design and Implementation of a Scalable Financial Exchange in the Public Cloud

### v8

@@ 4 Outbound Communication
On the exchange outbound side, market data is distributed to MPs fairly and with low latency. In on-prem exchanges, fairness—simultaneous data delivery to all MPs—is achieved by equalizing wire lengths between the exchange server and MPs and using low-jitter switches. Here, we develop overlay techniques atop a best-effort fabric to achieve fair and low-latency multicast that scales to a large number of receivers.
The simplest approach to transmit the same message to many receivers is multiple direct unicasts, where the sender sends the same message directly to each receiver. This increases the latency for later receivers: back-to-back unicasts are serialized over the sender’s egress, cumulatively increasing transmission and serialization delay.
Instead, we use an overlay tree to create a scalable multicast. Here, the sender sends messages to a few proxies, which relay them down the tree to more proxies, and recursively down to receiver VMs at the leaves. This proxy tree reduces the cumulative delay (Fig. 4) by limiting each node’s fanout. However, each additional tree hop introduces some latency. We minimize this by optimizing proxy packet processing with kernel bypass and zero-copy techniques.
@@
Additionally, tuning the tree’s fan-out () and depth () is crucial for minimizing latency for a given number of receivers (). We conducted experiments with various <> for , and and found that setting and achieves near-minimum latency. Given the cloud’s inherent variability, more sophisticated tuning does not significantly improve latency, as detailed in Appendix A.
@@
Additionally, we synchronize the clocks of all MPs, proxy nodes, and the ME using the Huygens clock synchronization algorithm [23]. As in CloudEx’s “hold-and-release” mechanism [4], we assign each multicast message a future timestamp called a deadline. After receiving a message, receivers hold it until its deadline before processing it to ensure fair dissemination of market data. The deadline is calculated to ensure all receivers likely receive a message before the deadline, and is set to the send time plus a conservative estimate of maximum latency from the sender to any of the receivers.
However, as the number of receivers increases and the latency shows high variance, the required holding duration by each receiver increases to ensure simultaneous delivery which inflates the multicast latency. When “hold-and-release” is used with our latency and latency variance reduction techniques, it scales gracefully as it only has to account for small residual latency variance thus requiring smaller holding duration to achieve fairness (§8.5).
We address the security issue of trusting the receivers to process messages at or after the deadline in §6. The implementation details of “hold-and-release” and the deadline calculation by scalably aggregating OWD estimates from all the receivers is presented in Appendix B. In the following, we describe techniques that enhance the proxy tree to reduce both latency and latency variance.
4.1 Round-Robin Packet Spraying
In round-robin packet spraying (RRPS), parent-child associations in the tree are flexible: proxies change their children after each multicast message. This provides two main benefits: (i) the effect of latency spikes is reduced and (ii) a sudden burst of market data is distributed among more network paths. Both benefits translate to reduced multicast latency.
@@ 4.1 Round-Robin Packet Spraying
In RRPS, for a given proxy, on every new multicast message, the set of children is circularly shifted by 1. For example, the first proxy in a layer will have the first proxies of the next layer as its children for message 0, the next proxies as children for message 1, etc. This ensures that each proxy has a parent proxy for each message, while the set of children for a parent continuously changes. With this technique, a single leaf node in a proxy tree has a total of different virtual paths starting from the root node. Successive messages traverse this abundance of virtual paths in a round-robin fashion. On the other hand, without RRPS, there is only one virtual path from the root to each leaf node and all messages traverse that path. The increased number of paths provides latency reduction benefits, as discussed below.
@@
First, if a spike occurs on any VM-to-VM virtual link, only a subset of messages is impacted: because messages are round-robined across all available virtual paths, many messages will take alternative paths that avoid this virtual link.222Here, we assume that spikes are characteristics of individual virtual links, not VMs themselves—as is often the case. If spikes were a VM characteristic—as is also the case sometimes—many virtual links would simultaneously be affected, reducing the benefits of RRPS. For such straggler VMs, we use proxy hedging (§4.2).
@@
Second, without RRPS, each VM communicates with a fixed set of children VMs for every multicast message. So a burst of messages has a higher chance of filling up queues in the network, inflating latency. However, RRPS uncovers and utilizes new virtual paths that were unused earlier. For example, consider two adjacent proxy layers: a parent layer with proxies and a children layer with proxies where each parent proxy has children. Without RRPS, the number of virtual paths utilized by messages going from parents to children is . With RRPS, there are virtual paths. This allows us to distribute bursts among many more virtual paths, which reduces queue build-up and latency.
@@
At its core, RRPS works because there is an abundance of virtual paths available within an overlay proxy tree. This is because every VM in a cloud region can communicate with every other VM since a VM-to-VM link is virtual—effectively the network topology is a clique. This is unlike multicast trees in physical networks where the network topology is not a clique and is limited by physical constraints like geographic distance and cables.
4.2 Proxy Hedging
@@
A VM’s performance in the public cloud can degrade due to factors like noisy neighbors or live VM migration, causing latency fluctuations for all messages passing through the VM. To mitigate this, we develop proxy hedging, where each node in the tree receives multiple copies of a message from proxies in the higher layer. The node processes only the first copy and discards duplicates, thus reducing the impact of any straggler proxy VM. This technique decreases both temporal and spatial latency variance, leading to more stable latency and reducing the delivery window size (the difference between the earliest and latest receiver’s latency). We first outline a basic design that offers latency improvements but results in high throughput wastage. Then, we introduce an optimized version that avoids excessive throughput wastage while maintaining latency benefits. For ease of explanation, we assume RRPS is not in effect, but the 2 techniques compose.
Strawman design: F+H Technique. Each (internal) tree node has extra children along with its original children, where is the hedging factor. For example, in a multicast tree with and (Figure 3), if we configure , we will add 1 hedge node in Layer-0 and 3 hedge nodes in Layer-1. Each hedge node will assist original proxies at the same layer and does not have children of its own. A hedge node assists these proxies by forwarding a message received from the hedge node’s parent to the hedge node’s nieces. In this way, a node receives copies of a message: 1 copy from its parent, and the other copies from the hedge nodes of the layer above it. This technique significantly reduces latency variance. However, it has a distinct drawback: Whereas a proxy only needs to multicast messages to the nodes, the hedge node needs to multicast the same message copies to nodes, creating an egress throughput bottleneck at the hedge node. We build on the observations to design the final scheme.
@@
Final design: SiblingHedging. Instead of adding extra hedge nodes in a tree, we let each proxy node also take on the responsibility of hedging. Each proxy node sends messages to the children of of its siblings along with sending messages to children of its own. For example, in Figure 3 (left), when hedging is not enabled (i.e., ), proxy ( proxy in layer) only receives messages from . As a result, may suffer from high latency if or the path from to encounters latency fluctuations. By using hedging, not only receives messages from , but also receives the same messages from one (if ) other node, . With proxy hedging, each node has more than one parent. This technique shows significant benefits that we evaluate in §8.2.
@@
With proxy hedging, a proxy processes the earliest received copy among copies of a message and discards the rest. This leads to significant latency reduction of each message, at the cost of some throughout wastage. We find out empirically as well as with Monte Carlo simulation that a small H (of up to 2) is enough for reaping most of the benefits of hedging. A Monte Carlo analysis of hedging is presented in appendix C that corroborates: (i) latency and its variance reduces as H increases, and (ii) increasing H shows diminishing returns.
@@
Proxy hedging leads to redundant work which decreases the effective throughput (i.e., goodput) of a proxy. The goodput of a proxy node in SiblingHedging can be defined as: because each proxy sends redundant messages to the children of ’s siblings. One way to recover the lost throughput is to employ several parallel trees that share the root and the leaves (assuming the leaves have enough ingress bandwidth). The increased number of proxies also raise concerns of dollar cost, which indeed is a drawback here. The dollar cost can be ameliorated by utilizing spot VMs for proxies instead of on-demand VMs as they are siginificantly cheaper [24]. However, utilizing spot VMs requires proper fault tolerance that we address in §7.
4.3 Receiver Hedging
If a VM belonging to an MP becomes a straggler, the MP would not be able to compete with others as the market data it receives will lag behind other MPs. To hedge this, we provide each MP with two receiver VMs. Each receiver VM belonging to an MP runs the same trading algorithm, receives market data, and generates orders. If any one VM belonging to an MP suffers from performance variation or data is not disseminated fairly to it, then the MP’s other VM would help retain good performance, with high probability. This technique improves the performance guarantees significantly at the tail.
Receiver hedging doubles the number of receiver VMs, which increases the dollar cost. On the other hand, due to the scaling properties of the proxy tree, the increase in OML (overall multicast latency) is negligible when the number of receiver VMs doubles. Moreover, when we consider latency for an MP as the minimum latency experienced by any of its two receiver VMs, the OML improves (§8.4). Appendix F discusses how the two VMs of an MP stay synchronized.
4.4 Remarks on Multicast Packet Losses

### v11

@@ 4. Market Data Multicast
To satisfy outbound fairness, market data should be distributed simultaneously to all MPs. We develop overlay techniques atop a best-effort fabric to achieve fair and low-latency multicast that scales to a large number of receivers.
Due to the lack of switch support, the multicast in the cloud can only be implemented by using multiple direct unicasts. Since the back-to-back unicasts are serialized over the sender’s egress, the latter receivers (among a large number of receivers) will receive the unicast message much later than the others due to the cumulative serialization delay at the sender. To tackle this unfairness issue, we choose to use an overlay tree to implement the scalable multicast. As illustrated in Figure 2, the sender sits at the root of the tree and only sends its messages to a limited number of proxies. Each proxy then relays the message down the tree to the lower-layer proxies, and recursively down to receiver VMs at the leaves. Since each node’s fanout is limited, the cumulative serialization delay is constrained at each layer, reducing the heterogeneity of message delays among receivers.
@@
Tuning the tree’s fan-out () and depth () is crucial for minimizing end-to-end multicast latency for a given number of receivers () as extra hops in the path of messages increase latency. We conducted experiments with various ¡¿ for , and and verified that and yield near-minimum latency, and more sophisticated strategies bring little gains beyond that (Appendix A) because of the cloud’s inherent variability.
@@
We synchronize the clocks of all MPs, and the ME using Huygens algorithm [Geng et al., 2018]. We follow CloudEx’s “hold-and-release” protocol [Ghalayini et al., 2021] and assign each multicast message a future timestamp called a deadline. After receiving a message, the gateways hold it until its deadline before forwarding to the MP’s trading algorithm. The deadline is calculated by adding a headroom to the sending time of the market data. To ensure all receivers likely receive a message before the deadline, the headroom is decided by the maximum latency from the sender to every receiver.
However, since Jasper targets a much larger number (~1000) of receivers than CloudEx(~10), we realize that simply applying the “hold-and-release” is not sufficient. As the number of receivers increases and the latency exhibits high variance, the required holding duration by a receiver increases to ensure simultaneous delivery to all receivers, which inflates the multicast latency. Tree helps in lowering the end-to-end latency, however latency variance still needs to be minimized. So, we incorporate three optimizations into our proxy tree to reduce both the end-to-end latency and the latency variance.
4.1. Round-Robin Packet Spraying
@@
In order to reduce latency spikes and alleviate the impact of bursts of market data, we develop round-robin packet spraying (RRPS). In RRPS, parent-child links in the tree are flexible, i.e., proxies change their children after each multicast message. Specifically, for a given proxy, on every new multicast message, the set of children is circularly shifted by 1. For example, the first proxy in a layer will have the first proxies of the next layer as its children for message 0, the next proxies as children for message 1, etc. This ensures that each proxy has a parent for each message, while the set of children for a parent proxy continuously changes. With RRPS, a single leaf node in a proxy tree has a total of different virtual paths starting from the root node. Successive messages traverse this abundance of virtual paths in a round-robin fashion. By contrast, without RRPS, there is only one virtual path from the root to each leaf node and all messages traverse that path. The increased path diversity provides two major benefits for latency reduction.
@@
First, if a spike occurs on any VM-to-VM virtual link, only a subset of messages is impacted: Since messages are round-robined across all available virtual paths, many messages will take alternative paths that avoid this virtual link.444Here, we assume that spikes are characteristics of individual virtual links, not VMs themselves—as is often the case. If spikes were a VM characteristic—as is also the case sometimes—many virtual links would simultaneously be affected, reducing the benefits of RRPS. For such straggler VMs, we use proxy hedging (§4.2).
@@
Second, RRPS uncovers and utilizes new virtual paths that were unused earlier. For example, consider two adjacent proxy layers: a parent layer with proxies and a children layer with proxies where each parent proxy has children. Without RRPS, the number of virtual paths utilized by messages going from parents to children is . With RRPS, there are virtual paths. This allows us to distribute bursts among many more virtual paths, which reduces queue build-up and latency.
@@
At its core, RRPS works because there is an abundance of virtual paths available within an overlay proxy tree. This is because every VM in a cloud region can communicate with every other VM since a VM-to-VM link is virtual—effectively the network topology is a clique. This is unlike the links in physical networks [Andersen et al., 2003] where the network topology is not a clique and is limited by physical constraints like geographic distance and cables.
4.2. Proxy Hedging
@@
A VM’s performance in the cloud can degrade due to factors like noisy neighbors or live VM migration, causing latency fluctuations for all messages passing through the VM [Sachidananda, ]. To mitigate this, we develop proxy hedging, where each node in the tree receives multiple copies of a message from proxies in the higher layer. A node processes only the first copy and discards duplicates, thus reducing the impact of any straggler proxy VMs. It decreases both temporal and spatial latency variance, leading to more stable latency and reducing the delivery window size (i.e., the difference between the earliest and latest receiver’s latency). For ease of explanation, we assume RRPS is not in effect, but the 2 techniques compose.
@@
In proxy hedging, each proxy node sends messages to the children of of its siblings along with sending messages to children of its own where is the hedging factor. For example, in Figure 2 (left), when hedging is not enabled (i.e., ), proxy ( proxy in layer) only receives messages from . As a result, may suffer from high latency if or the path from to encounters latency fluctuations. By using hedging, not only receives messages from , but also receives the same messages from one (if ) other node, . This technique shows significant benefits that we evaluate in §6.5.1.
@@
A proxy processes the earliest received copy among copies of a message and discards the rest, achieving significant latency reduction of each message, at the cost of some throughput wastage. We find out empirically as well as with Monte Carlo simulation that a small (of up to 2) is sufficient for reaping its most benefits. A Monte Carlo analysis of hedging is presented in Appendix C that corroborates: (i) latency and its variance reduces as increases, and (ii) increasing shows diminishing returns.
@@
Unlike the previous technique, RRPS, proxy hedging leads to redundant work which decreases the effective throughput (i.e., goodput) of a proxy. The goodput of a proxy node can be defined as: because each proxy sends redundant messages to the children of siblings. One way to recover the lost throughput is to employ several parallel trees that share the root and the leaves (assuming the leaves have enough ingress bandwidth). The increased number of proxies also raises concerns of dollar cost, which could potentially be a drawback here. Appendix H includes the detailed estimation of monetary cost.
4.3. Receiver Hedging
If a VM belonging to an MP becomes a straggler, the MP would become more likely to lose the trading competition against the other MPs, because the market data it receives will lag behind other MPs. To hedge this, we provide each MP with two receiver VMs; similar measures are taken in on-premises exchanges so we do not expand more on it. This simple technique enhances performance significantly at the tail, more details in Appendix E.
4.4. Remarks on Multicast Packet Losses